OpenStudy (anonymous):

between what two consecutive integers does square root of 379 lie?

OpenStudy (anonymous):

say x is one number, next number (consecutive) is x+1 Product of these two will give 379\[x(x+1)=379 \rightarrow x^2 + x =379 \rightarrow x^2+x-379 =0\] Solve fo x

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=complete%20the%20square%20x^2%2Bx-379&lk=2 roots are approx 19, 20. and, yes, they are the consecutive numbers.